OpenStudy (qwertykingghlpo):

Algebra help.... How many real number solutions does the equation have? 0=-4x^2+7x-8 A. no solutions B. one solution C. two solutions D. Infinitely many solutions

ganeshie8 (ganeshie8):

find the discriminant

ganeshie8 (ganeshie8):

\(-4x^2+7x-8 ~~=~~0\) compare this with the standard equation of parabola \(ax^2+bx+c~~=~~0\)

ganeshie8 (ganeshie8):

\(a=?\) \(b=?\) \(c=?\)

ganeshie8 (ganeshie8):

I am doing that. Could you answer my previous q

ganeshie8 (ganeshie8):

have you understood my q ?

ganeshie8 (ganeshie8):

like, what exactly im asking

ganeshie8 (ganeshie8):

please scroll up and try to attempt my question again

ganeshie8 (ganeshie8):

this one

ganeshie8 (ganeshie8):

Not quite, looks you forgot the signs

ganeshie8 (ganeshie8):

try again

ganeshie8 (ganeshie8):

good, next work the discriminant

ganeshie8 (ganeshie8):

discriminant = b^2 - 4ac

ganeshie8 (ganeshie8):

just plugin the given values and simplify

ganeshie8 (ganeshie8):

take ur time

ganeshie8 (ganeshie8):

Yes simplify